integrate the following integral
∫ sin^2(x/2) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int sin^{2}\frac{x}{2}dx
I = \frac{1}{2}\int (1-cosx)dx
I = \frac{1}{2}[\int dx-\int cosxdx]
I = \frac{1}{2}[x-\int cosxdx]
I = \frac{1}{2}[x-sinx]+c